OpenStudy (anonymous):

What is the ordered pair that represents the solution of this system of equations. y = - x + 6 and y = x – 2.

OpenStudy (anonymous):

I'd put both equations equal to each other. -x+6 = x-2 -2x = -8 x = 4 To find y, substitute 4 back into the equations: -4 + 6 = 2 4 - 2 = 2 the y value is 2 (4,2)

OpenStudy (anonymous):

Thanks! :)

OpenStudy (pratyush5):

We can also do it by adding both equations.